# Simultaneous Proton and Electron Energization during Macroscale Magnetic Reconnection

### v1

@@ 3 Simulation Results
Figure 2 shows the energy per particle (energy density divided by number density) of the particle protons and particle electrons, $\langle W\rangle$, and the firehose parameter, $\alpha$ = $1-4\pi(P_{\parallel}-P_{\perp})/B^{2}$, at late time, $t/\tau_{A}=13$, from the same simulation. In the top two panels darker colors represent higher average energies per particle. Thus, the average energy of the protons is significantly greater than that of the electrons. In the bottom panel the firehose parameter is shown. Notably, extensive areas within the magnetic islands approach marginal stability ($\alpha=0$), punctuated by small unstable regions (red shades). Consequently, the local magnetic tension, a crucial factor in driving reconnection and particle energy gain, is strongly reduced within the large flux rope and somewhat reduced in the smaller flux ropes. This highlights the critical importance of particle feedback on the MHD fluid dynamics. Traditional MHD models that include test-particle dynamics do not include this feedback and thereby risk an uncontrolled increase of the particle energy.
@@
At late time, after a single large magnetic island dominates the current sheet, we calculate the proton energy spectra by aggregating particle counts throughout the entire simulation domain. Figure 3(a) displays the differential number densities $F(W)=dN(W)/dW$ plotted against normalized energy $W/m_{i}C_{A0}^{2}$ at various times for $B_{g}/B_{0}=0.25$. $F(W)$ begins as a Maxwellian but develops a power-law tail, characterized by an index $\delta^{\prime}$, that first appears for low energies before eventually extending to near the maximum energy observed in the domain. The powerlaw index at late time approaches 3.4. Similar behavior was observed for electrons in the earlier electron-only version of kglobal (Arnold et al., 2021).
@@
Figure 4: The late-time distributions of energetic protons in a simulation with $B_{g}/B_{0}=0.25$ for various values of $S_{\nu}$ (effective system size) show that the power-law indices of the spectra for the different ratios are nearly the same. The only notable difference is that as the size of the system increases, the maximum energy of the protons also increases.
In addition, we explore the late-time distributions of energetic protons for a guide field $B_{g}/B_{0}=0.25$ and various values of $S_{\nu}$ (effective system size). The results in Fig. 4 show that the power-law indices of the spectra become independent of the system size for large domains. However, the total energy contained in the non-thermal protons increases with reconnected flux with the result that larger systems produce power laws that extend to higher energy.
@@
To compare our results to observations, consider that in solar flare current sheets the typical magnetic field magnitude is about 50 $\mathrm{Gauss}$ and the density of protons is around $10^{10}~{}\mathrm{cm}^{-3}$, which corresponds to an Alfvén speed of 1000 km/s. For our simulation, we initially set the temperatures of protons and electrons to be $0.0625m_{i}C_{A0}^{2}$. Given that $m_{i}C_{A0}^{2}$ is approximately 10 keV in the solar case, this translates to an initial plasma temperature in the simulations of around $0.6$ keV, which exceeds typical ambient coronal values in the absence of prior reconnection energy release. On the other hand, the late time energy spectra of both species extend to far larger energies, so the initial temperatures are not expected to significantly influence these late time spectra. The results shown in Figure 6 indicate that the maximum energy attained by energetic electrons can exceed $100m_{i}C_{A}^{2}\sim 1$ MeV, while energetic protons can achieve energies in excess of $500m_{i}C_{A}^{2}\sim 5$ MeV. These values are consistent with findings from gamma-ray flares documented by RHESSI (Lin et al., 2003). A similar scaling of the results from the largest simulation domain (see Figure 4) suggests that the proton energy can exceed 20 MeV.
@@
In Figure 7 we show $F(W)$ (black) for protons and electrons at late time after the distribution functions have reached a steady state. The data is taken from within the large late-time island shown in Figure 1(e). Thus, the spectra shown include only particles that have undergone heating and energization during reconnection. The spectra are fit (represented in solid red) by the sum of a Maxwellian distribution and a kappa distribution (Arnold et al., 2021). This procedure is designed to capture the characteristics of both hot thermal and nonthermal particle populations. From this fitting approach, we extract several parameters:
@@
Type Guide Field Total Maxwellian Kappa Kappa Nonthermal Powerlaw
Maxwellian Nonthermal Break Point Index
(in $B_{0}$) (in $m_{i}C_{A0}^{2}$) (in $m_{i}C_{A0}^{2}$) (in $m_{i}C_{A0}^{2}$) (in $m_{i}C_{A0}^{2}$) (in $m_{i}C_{A0}^{2}$) ($\delta^{\prime}$)
Electron 0.25 0.202 0.059 0.048 0.095 0.28 3.3
Electron 0.50 0.158 0.101 0.033 0.024 0.35 5.4
Electron 0.75 0.135 0.113 0.016 0.006 0.50 8.8
Proton 0.25 0.581 0.050 0.191 0.340 1.34 3.4
Proton 0.50 0.295 0.035 0.141 0.119 0.72 4.9
Proton 0.75 0.200 0.051 0.104 0.045 0.86 7.7
@@
Table 3 shows the relative energy distribution between thermal and non-thermal electrons and protons at late time, again as a function of guide field strength. In this data the total number density of particles is normalized to unity and the numbers in the table reflect the energy content of the various components in units of $m_{i}C_{A}^{2}$. As the guide field increases, the energy in the non-thermal component of the kappa distribution decreases for both electrons and protons. For electrons, the kappa non-thermal energy decreases from 0.095 at a guide field of 0.25 to 0.006 at a guide field of 0.75. For protons, the kappa non-thermal energy decreases from 0.340 at a guide field of 0.25 to 0.045 at a guide field of 0.75. Additionally, the break point energy, which marks the transition from thermal to non-thermal particle dominance in the spectra, behaves differently for electrons and protons. For electrons, as the guide field increases, the break point shifts to higher values, while for protons the break point first decreases with increasing guide field and then the trend reverses. The powerlaw index also increases with the guide field, reflecting a steep drop-off in the number of non-thermal particles at higher energies.
@@ 4 Conclusion
In this paper, we present the results of a numerical simulation that produces the self-consistent and simultaneous energization of electrons and protons to non-thermal energies due to magnetic reconnection in a macroscale system. We calculate the time evolution of macroscale systems, starting from a thin current sheet, progressing through many reconnection events, and culminating in the formation of a large merged island at a late time. Electrons and protons gain energy through Fermi reflection as multiple flux ropes grow and merge. At late time the energetic particles have mostly accumulated into the large merged island as seen in earlier electron-only simulations (Arnold et al., 2021). We also present an example of the firehose stability parameter at late time. The core of magnetic flux ropes approach firehose marginal stability so that the magnetic tension within flux ropes (the drive mechanism for reconnection) is suppressed. It is through the firehose parameter that energetic protons and electrons feed back onto the MHD fluid.
@@
The proton and electron spectra both form power-law distributions at high energy with comparable spectral indices. As the magnitude of the guide field decreases, both protons and electrons exhibit a harder energy spectra so that there are larger numbers of high-energy particles. Similar trends have been previously reported (Dahlin et al., 2014; Arnold et al., 2021; Zhang et al., 2021). The power-law slopes are generally insensitive to the system size. However, the total energy in non-thermal particles increases with reconnected flux, especially in larger systems. The powerlaw spectra of both species extend nearly three decades in energy for large domains. Our results also indicate a distinct difference in energy gain between protons and electrons: protons gain more energy and their maximum energy exceeds that of electrons.
Energy is normalized to $m_{i}C_{A}^{2}$ in our simulations. A typical value of this parameter in a solar flare is 10 keV (see Sec. 3). Thus, our simulations suggest that energetic electrons can reach energies as high as $10^{2}m_{i}C_{A}^{2}\sim 1$MeV, while energetic protons can reach beyond $10^{3}m_{i}C_{A}^{2}\sim 10$ MeV The simulations further reveal that the break point, which is the transition from thermal to non-thermal particle dominance in the particle spectra, is around $1.34m_{i}C_{A}^{2}\sim 13$keV for protons and $0.28m_{i}C_{A}^{2}\sim 3$ keV for electrons. The powerlaw spectra of protons should extend down to far lower energies than the typical 1 MeV minimum proton energy that is measureable in flares (Lin et al., 2003; Emslie et al., 2012). Of course, the powerlaw spectra of a variety of ion species from impulsive flares have been documented in spacecraft measurements in the solar wind (Reames, 1999; Mason, 2007). This suggests that the total energy content of protons in solar flares is likely to be greater than previously expected (Emslie et al., 2012).
@@
Magnetic reconnection in low density events in the Earth’s magnetotail are associated with high values of $m_{i}C_{A}^{2}$ with Alfvén speeds reaching values above 1000km/s. Powerlaw distributions of electrons (Øieroset et al., 2002; Ergun et al., 2020) as well as protons (Ergun et al., 2020) in such events have been documented. In situ measurements of reconnection outflows have documented that electron (Phan et al., 2013) and proton heating (Drake et al., 2009; Phan et al., 2014) scale with $m_{i}C_{A}^{2}$ and recently the observations have been found to extend to the energetic tail events, with electron temperature increments reaching several keV (Øieroset et al., 2023). Simulations carried out with particle-in-cell codes also support this scaling of particle energy gain during reconnection (Haggerty et al., 2015). Since the energy normalization parameter in our simulations is $m_{i}C_{A}^{2}$, the results of our simulations shown in Table 3 also support this scaling. The in situ observations have established that proton bulk heating during reconnection is substantially greater than that of electrons by a factor of around 6. The data in Table 3 indicates that the ratio of proton to electron total energy is around 3 for the smallest guide field simulation (the observations are mostly limited to weak guide field events), somewhat less than in the observations. On the other hand, in the Earth space environment the protons typically have initial temperatures well above that of electrons. This might account for their greater energy gain since the Fermi heating and acceleration is strongest for particles with greater energy.
@@
Thus, simulations with the kglobal model can be carried out to explore reconnection-driven heating and energization in realistic macroscale systems. By using measured values of $m_{i}C_{A}^{2}$, the initial temperatures of protons and electrons, and ambient guide fields, the simulation model can produce results for direct comparison with observational data from solar flares and in situ measurements from the Earth space environment and and the solar wind.

### v2

@@ 3 Simulation Results
Figure 2 shows the energy per particle (energy density divided by number density) of the particle protons and particle electrons, $\langle W\rangle$, and the firehose parameter, $\alpha$ = $1-4\pi(P_{\parallel}-P_{\perp})/B^{2}$, at late time, $t/\tau_{A}=13$, from the same simulation. In the top two panels darker colors represent higher average energies per particle. Thus, the average energy of the protons is significantly greater than that of the electrons.
The firehose parameter in the bottom panel of Figure 2 is a measure of the impact of pressure anisotropy on the local magnetic tension, given by $\alpha\mathbf{B}\cdot\mathbf{\nabla}\mathbf{B}/4\pi$. Specifically, when $\alpha$ approaches zero the magnetic tension, which is the driver of magnetic reconnection also goes to zero. Thus, the dominant feedback mechanism of energetic particles on reconnection dynamics is through the reduction of magnetic tension as measured by the firehose parameter. In the context of the simulation from kglobal, both $P_{\parallel}$ and $P_{\perp}$ are derived from the particle species since both the fluid proton and electrons are taken to be isotropic and don’t impact the magnetic tension force. Thus, in Figure 2 extensive areas within the magnetic islands approach marginal stability ($\alpha=0$), punctuated by small unstable regions (red shades). Consequently, the local magnetic tension, a crucial factor in driving reconnection and particle energy gain, is strongly reduced within the large flux rope and somewhat reduced in the smaller flux ropes. This highlights the critical importance of particle feedback on the MHD fluid dynamics. Traditional MHD models that include test-particle dynamics do not include this feedback and thereby risk an uncontrolled increase of the particle energy.
@@
At late time, after a single large magnetic island dominates the current sheet, we calculate the proton energy spectra by aggregating particle counts throughout the entire simulation domain. Figure 3(a) displays the differential number densities $F(W)=dN(W)/dW$ plotted against normalized energy $W/m_{i}C_{A0}^{2}$ at various times for $B_{g}/B_{0}=0.25$. $F(W)$ begins as a Maxwellian but develops a power-law tail, characterized by an index $\delta^{\prime}$, that first appears for low energies before eventually extending to near the maximum energy observed in the domain. The powerlaw index at late time approaches 3.4. Similar behavior was observed for electrons in the earlier electron-only version of kglobal (Arnold et al., 2021). Note that the particle distribution at low energy in Figure 3(a) does not change significantly in time. This is not because the protons are not heated during reconnection but because the distribution in the figure is from the entire computational domain, which includes large regions of plasma upstream of the reconnecting current layer that has not been accelerated. The heating of the plasma will be more evident when we focus on the particle distributions within the large islands that develop at late time in the simulations.
@@
Figure 4: The late-time distributions of energetic protons in a simulation with $B_{g}/B_{0}=0.25$ for various values of $S_{\nu}$ (effective system size) reveal that the power-law indices of the spectra vary weakly with the domain size. The rightmost point of each curve marks the one-count level, indicating the flux when there is just one particle count per energy bin. The notable difference is that as the size of the system increases, the maximum energy of the protons also increases.
In addition, we explore the late-time distributions of energetic protons for a guide field $B_{g}/B_{0}=0.25$ and various values of $S_{\nu}$ (effective system size). The results in Fig. 4 show that the power-law indices of the spectra are only weakly dependent on the system size for large domains. As the size of the system increases, the spectra become slightly harder and the upper limit in the particle energy gain increases. This is because in larger domains more flux ropes form in the reconnecting current sheet and it is the merger of those flux ropes that efficiently drives particle energy gain.
@@
To compare our results to observations, consider that in solar flare current sheets the typical magnetic field magnitude is about 50 $\mathrm{Gauss}$ and the density of protons is around $10^{10}~{}\mathrm{cm}^{-3}$, which corresponds to an Alfvén speed of 1000 km/s. For our simulation, we initially set the temperatures of protons and electrons to be $0.0625m_{i}C_{A0}^{2}$. Given that $m_{i}C_{A0}^{2}$ is approximately 10 keV in the solar case, this translates to an initial plasma temperature in the simulations of around $0.6$ keV (around 7 MK), which exceeds typical ambient coronal values in the absence of prior reconnection energy release. On the other hand, the late time energy spectra of both species extend to far larger energies, so the initial temperatures are not expected to significantly influence these late time spectra. The results shown in Figure 6 indicate that the maximum energy attained by energetic electrons can exceed $100m_{i}C_{A}^{2}\sim 1$ MeV, while energetic protons can achieve energies in excess of $500m_{i}C_{A}^{2}\sim 5$ MeV. These values are consistent with findings from gamma-ray flares documented by RHESSI (Lin et al., 2003). A similar scaling of the results from the largest simulation domain (see Figure 4) suggests that the proton energy can exceed 20 MeV.
@@
In Figure 7 we show $F(W)$ (black) for protons and electrons at late time after the distribution functions have reached a steady state. The data is taken from within the large late-time island shown in Figure 1(e). Thus, the spectra shown include only particles that have undergone heating and energization during reconnection. Shown in Figure 8 are the intial and final electron and proton distributions from the same large island but using a linear scale. This figure more clearly shows the changes in the distribution of low energy particles as they are heated during reconnection. The spectra in Figure 7 are fit (solid red) by the sum of a Maxwellian distribution and a kappa distribution (Arnold et al., 2021). This procedure is designed to capture the characteristics of both hot thermal and nonthermal particle populations. From this fitting approach, we extract several parameters:
@@
Figure 8: The comparison between the initial spectra and late-time spectra of protons and electrons is presented on a linear-linear scale to more clearly reveal the changes in the particle distributions at low energy.
@@
Type Guide Total Maxwellian Kappa Kappa Thermal Nonthermal Powerlaw
Field Maxwellian Nonthermal Temperature Break Point Index
(in $B_{0}$) (in $m_{i}C_{A0}^{2}$) (in $m_{i}C_{A0}^{2}$) (in $m_{i}C_{A0}^{2}$) (in $m_{i}C_{A0}^{2}$) (in $m_{i}C_{A0}^{2}$) (in $m_{i}C_{A0}^{2}$) ($\delta^{\prime}$)
Electron 0.25 0.202 0.059 0.048 0.095 0.086 0.28 3.3
Electron 0.50 0.158 0.101 0.033 0.024 0.094 0.35 5.4
Electron 0.75 0.135 0.113 0.016 0.006 0.087 0.50 8.8
Proton 0.25 0.581 0.050 0.191 0.340 0.184 1.34 3.4
Proton 0.50 0.295 0.035 0.141 0.119 0.132 0.72 4.9
Proton 0.75 0.200 0.051 0.104 0.045 0.109 0.86 7.7
@@
Table 3 shows the relative energy of thermal and non-thermal electrons and protons at late time, again as a function of guide field strength. In this data the total number density of particles is normalized to unity and the numbers in the table reflect the energy content of the various components in units of $m_{i}C_{A}^{2}$. As the guide field increases, the energy in the non-thermal component of the kappa distribution decreases for both electrons and protons. For electrons, the kappa non-thermal energy decreases from 0.095 at a guide field of 0.25 to 0.006 at a guide field of 0.75. For protons, the kappa non-thermal energy decreases from 0.340 at a guide field of 0.25 to 0.045 at a guide field of 0.75. The downstream thermal temperatures are also shown. For electrons the thermal temperature, represented as a weighted sum of the Maxwellian and Kappa Maxwellian distributions, remains comparable for different guide fields. The temperature increase is modest compared with the initial value of $0.0625m_{i}C_{A0}^{2}$. The heating is driven by the macroscale parallel electric field as explored by Arnold et al. (2021). In contrast, protons exhibit a much larger increase in temperature as has been documented in the Earth space environment (Phan et al., 2013b, 2014; Øieroset et al., 2023). For the lowest guide field in Table 3 the ratio of the temperature increment of the protons to electrons is around five, which is slightly smaller than in measurements in the Earth space environment (Phan et al., 2014). Also noticeable is the decrease in thermal temperature of protons with increasing guide field. We emphasize that because the data in Table 3 is normalized to $m_{i}C_{A0}^{2}$, the temperature increments of both species scales with this parameter as in the observations (Drake et al., 2009; Phan et al., 2013b, 2014; Øieroset et al., 2023).
Also shown in Table 3 is the breakpoint energy, which marks the transition from thermal to non-thermal particle dominance in the spectra. The depedence of the breakpoint energy on the guide field behaves differently for electrons and protons. For electrons, as the guide field increases, the break point shifts to higher values, while for protons the break point first decreases with increasing guide field and then the trend reverses. The powerlaw index also increases with the guide field, reflecting a steep drop-off in the number of non-thermal particles at higher energies.
@@ 4 Conclusion
In this paper, we present the results of numerical simulations of the self-consistent energization of electrons and protons during magnetic reconnection in a macroscale system. We calculate the time evolution of the macroscale system, starting from a thin current sheet, progressing through the formation of multiple flux ropes and their merger, culminating in the formation of a large merged island at late time. Electrons and protons gain energy through Fermi reflection as multiple flux ropes grow and merge. At late time the energetic particles have mostly accumulated into the large merged island as seen in earlier electron-only simulations (Arnold et al., 2021). We also present an example of the firehose stability parameter at late time. The core of magnetic flux ropes approach firehose marginal stability so that the magnetic tension within flux ropes (the drive mechanism for reconnection) is suppressed. It is through the firehose parameter that energetic protons and electrons feed back onto the MHD fluid during magnetic reconnection.
@@
The proton and electron spectra both form power-law distributions at high energy with comparable spectral indices. As the magnitude of the guide field decreases, both protons and electrons exhibit a harder energy spectra so that there are larger numbers of high-energy particles. Similar trends have been previously reported (Dahlin et al., 2014; Arnold et al., 2021; Zhang et al., 2021). The power-law slopes are generally insensitive to the system size. However, the highest energy of non-thermal particles increases in larger systems, which reflects the larger number of flux ropes undergoing merger to drive particle energy gain. The powerlaw spectra of both species extend nearly three decades in energy for large domains. Our results also indicate a distinct difference in energy gain between protons and electrons: protons gain more energy and their maximum energy exceeds that of electrons.
[…]
@@
Magnetic reconnection in low density events in the Earth’s magnetotail are associated with high values of $m_{i}C_{A}^{2}$ with Alfvén speeds reaching values above 1000km/s. Powerlaw distributions of electrons (Øieroset et al., 2002; Ergun et al., 2020) as well as protons (Ergun et al., 2020) in such events have been documented. In situ measurements of reconnection outflows have documented that electron (Phan et al., 2013a) and proton heating (Drake et al., 2009; Phan et al., 2014) scale with $m_{i}C_{A}^{2}$ and recently the observations have been found to extend to the energetic tail events, with electron temperature increments reaching several keV (Øieroset et al., 2023). Simulations carried out with particle-in-cell codes also support this scaling of particle energy gain during reconnection (Haggerty et al., 2015). Since the energy normalization parameter in our simulations is $m_{i}C_{A}^{2}$, the results of our simulations shown in Table 3 also support this scaling. The in situ observations have established that proton bulk heating during reconnection is substantially greater than that of electrons by a factor of around 8. The data in Table 3 indicates that the ratio of the increments of proton to electron temperature is around 5 for the smallest guide field simulation (the observations are mostly limited to weak guide field events), somewhat less than in the observations. On the other hand, in the Earth space environment the protons typically have initial temperatures well above that of electrons. This might account for their greater energy gain since the Fermi heating and acceleration is strongest for particles with greater energy.
@@
Thus, simulations with the kglobal model can be carried out to explore reconnection-driven heating and energization in macroscale systems. By using measured values of $m_{i}C_{A}^{2}$, the initial temperatures of protons and electrons, and ambient guide fields, the simulation model can produce results for direct comparison with observational data from solar flares and in situ measurements from the Earth space environment and and the solar wind.